# MTMamba++: Enhancing Multi-Task Dense Scene Understanding via Mamba-Based Decoders

### v1

@@ 4.3 Model Analysis
To analyze the proposed MTMamba++, we first introduce two baselines: (i) “Multi-task” represents an MTL model using two standard Swin Transformer blocks [30] after the ECR block in each decoder stage for each task; and (ii) “Single-task” is the single-task counterpart of “Multi-task” (i.e., each task has a task-specific encoder-decoder). Without specific instructions, the encoder in this section is the Swin-Large Transformer.
@@ 4.3.1 Effectiveness of Each Component
The decoders of MTMamba++ contain two types of core blocks: STM and CTM blocks. Compared to the preliminary version MTMamba [29], MTMamba++ replaces the F-CTM block and DenseHead of MTMamba with the S-CTM block and LiteHead, respectively. In this experiment, we study the effectiveness of each component on the NYUDv2 dataset. The results are shown in Table III, where #5 and #7 are the default configurations of MTMamba and MTMamba++, respectively.
Firstly, the STM block outperforms the Swin Transformer block [30] in terms of efficiency and effectiveness for multi-task dense prediction, as indicated by the superior results in Table III (#2 vs. #3). Secondly, merely increasing the number of STM blocks from two to three does not enhance performance significantly. When the F-CTM block is incorporated, the performance largely improves in terms of $\Delta_{m}$ (#5 vs. #3/#4), demonstrating the effectiveness of F-CTM. Thirdly, the comparison between #5 and #6 shows that LiteHead is more effective and efficient than DenseHead. Fourthly, compared #6 with #7, we can find that replacing F-CTM with S-CTM leads to a significant performance improvement in all tasks with a tiny additional cost, demonstrating that the semantic-aware interaction in S-CTM is more effective than F-CTM. Finally, the default configuration of MTMamba++ (i.e., #7) significantly surpasses the “Single-task” approach across all tasks, thereby demonstrating the effectiveness of MTMamba++.
TABLE IV: Comparison between SSM and attention on NYUDv2. We replace the SSM-related modules in MTMamba++ (i.e., the SS2D and CSS2D modules) with attention-based mechanisms.
@@
SS2D $\rightarrow$ self-attention & CSS2D $\rightarrow$ cross-attention 55.15 0.4945 18.72 79.00 +2.63 448 796
MTMamba++ 57.01 0.4818 18.27 79.40 +4.82 315 524
@@
TABLE VII: Effect of expand factor $\alpha$ in MTMamba++ on NYUDv2. $\alpha=2$ is the default configuration.
$\alpha$ Semseg Depth Normal Boundary $\Delta_{m}$[%] #Param FLOPs
@@
$1$ 56.74 0.4927 18.45 79.40 +3.93 276 395
$2$ 57.01 0.4818 18.27 79.40 +4.82 315 524
$3$ 55.85 0.4882 18.40 79.10 +3.71 354 653
4.3.2 Comparison between SSM and Attention
To demonstrate the superiority of the SSM-based architecture in multi-task dense prediction, we conduct an experiment on NYUDv2 by replacing the SSM-related operations in MTMamba++ with attention-based mechanisms. Specifically, the SS2D module in the STM block is replaced with the window-based multi-head self-attention [30], while the CSS2D module in the S-CTM block is substituted with the window-based multi-head cross-attention. The results are shown in Table IV. As can be seen, the SSM-based method (i.e., MTMamba++) is more effective and efficient than the attention-based method, which demonstrates that SSM is more powerful in dealing with multi-task dense prediction tasks.
4.3.3 Effectiveness of Each Decoder Stage
As shown in Figure 1, the decoder of MTMamba++ consists of three stages. In this experiment, we study the effectiveness of these three stages on the NYUDv2 dataset. Table V shows the results, which clearly demonstrate that each stage contributes positively to the performance of MTMamba++.
@@ 4.3.4 Effect of Each Scan in CSS2D Module
4.3.5 Effect of $\alpha$
As mentioned in Sections 3.5 and 3.6.2, in MTMamba++, both STM and S-CTM blocks expand the feature channel to improve the model capacity by a hyperparameter $\alpha$. We conduct an experiment on NYUDv2 to study the effect of $\alpha$. The results are shown in Table VII. As can be seen, increasing $\alpha$ leads to a significant increase in model size and FLOPS cost. As for performance, $\alpha=2$ performs the best and is thus used as the default configuration in MTMamba++.
@@ 4.3.6 Performance on Different Encoders
Figure 5: Qualitative comparison with state-of-the-art methods (i.e., InvPT [7], TaskPrompter [10], and MTMamba [29]) on the NYUDv2 dataset. MTMamba++ generates better predictions with more accurate details as marked in yellow circles. In the semantic segmentation task, the black regions in GT denote the background and are excluded from the computation of loss and evaluation metric (i.e., mIoU). Zoom in for more details.
Figure 6: Qualitative comparison with state-of-the-art methods (i.e., InvPT [7], TaskPrompter [10], and MTMamba [29]) on the PASCAL-Context dataset. MTMamba++ generates better predictions with greater precision in the details, as marked in yellow circles. Zoom in for more details.
Figure 7: Qualitative comparison with state-of-the-art methods (i.e., InvPT [7], TaskPrompter [10], and MTMamba [29]) on the Cityscapes dataset. MTMamba++ produces better predictions with more precise details. Zoom in for more details.
@@ 4.3.7 Analysis of Model Size and Cost
In Table IX, we compare the model size and FLOPs between the proposed MTMamba++ and baselines on the PASCAL-Context dataset. As can be seen, with comparable parameter size and computational cost, MTMamba++ significantly outperforms previous methods on four of five tasks.
@@ 4.4 Visualization of Predictions
In this section, we compare the output predictions from MTMamba++ against the state-of-the-art methods, including InvPT [7], TaskPrompter [10], and MTMamba [29]. Figures 5, 6, and 7 show the qualitative results on NYUDv2, PASCAL-Context, and Cityscapes datasets, respectively. As can be seen, MTMamba++ has better visual results than baselines in all datasets. For example, as highlighted with yellow circles in Figure 5, MTMamba++ generates more accurate results with better alignments for the semantic segmentation task. Figure 6 demonstrates that MTMamba++ produces better predictions with more accurate details (as highlighted in yellow circles) for both semantic segmentation and human parsing tasks and more distinct boundaries for the object boundary detection task. Similarly, Figure 7 illustrates that MTMamba++ delivers more precise segmentation with superior alignment. Hence, both qualitative study (Figures 5, 6, and 7) and quantitative study (Tables I and II) show the superior performance of MTMamba++.
@@ 5 Conclusion
In this paper, we propose MTMamba++, a novel multi-task architecture with a Mamba-based decoder for multi-task dense scene understanding. With two types of core blocks (i.e., STM and CTM blocks), MTMamba++ can effectively model long-range dependency and achieve cross-task interaction. We design two variants of the CTM block to promote knowledge exchange across tasks from the feature and semantic perspectives, respectively. Experiments on three benchmark datasets demonstrate that the proposed MTMamba++ achieves better performance than previous CNN-based and Transformer-based methods.

### v2

@@ 4.3 Model Analysis
In this section, we provide a comprehensive analysis of the proposed MTMamba++. Without specific instructions, the encoder in this section is the Swin-Large Transformer.
@@ 4.3.1 Effectiveness of Each Component
The decoders of MTMamba++ contain two types of core blocks: STM and CTM blocks. Compared to the preliminary version MTMamba [24], MTMamba++ replaces the F-CTM block and DenseHead of MTMamba with the S-CTM block and LiteHead, respectively.
In this experiment, we study the effectiveness of each component on the NYUDv2 dataset. We first introduce two groups of baselines: (i) “Multi-task” represents an MTL model using only standard Swin Transformer blocks [25] after the ECR block in each decoder stage for each task; and (ii) “Single-task” means that each task has a task-specific encoder-decoder. The results are shown in Table III, where #9 and #11 are the default configurations of MTMamba and MTMamba++, respectively.
Firstly, the STM block outperforms the Swin Transformer block [25] in terms of efficiency and effectiveness for multi-task dense prediction, as indicated by the superior results in Table III (#3 vs. #7 and #5 vs. #8). Secondly, merely increasing the number of STM blocks from two to three does not enhance performance significantly. When the F-CTM block is incorporated, the performance largely improves in terms of $\Delta_{m}$ (#9 vs. #7/#8), demonstrating the effectiveness of F-CTM. Thirdly, comparisons between #3 and #4, #5 and #6, as well as #9 and #10 show that LiteHead is more effective and efficient than DenseHead. Fourthly, compared #10 with #11, we can find that replacing F-CTM with S-CTM leads to a significant performance improvement in all tasks with a tiny additional cost, demonstrating that the semantic-aware interaction in S-CTM is more effective than F-CTM. Finally, the default configuration of MTMamba++ significantly surpasses the “Single-task” baselines across all tasks (#11 vs. #1/#2), thereby demonstrating the effectiveness of MTMamba++.
4.3.2 Comparison between SSM and Attention
To demonstrate the superiority of the SSM-based architecture in multi-task dense prediction, we conduct an experiment on NYUDv2 by replacing the SSM-related components in MTMamba++ with attention-based counterparts. Specifically, we substitute the SS2D module in the STM block with window-based multi-head self-attention [25] and replace the CSS2D module in the S-CTM block with window-based multi-head cross-attention. The comparative results in Table IV show that MTMamba++ significantly outperforms the attention-based variant across all tasks while requiring approximately 29.7% fewer parameters and 34.2% lower FLOPs. This efficiency advantage is primarily from SSM’s linear computational complexity with respect to sequence length, in contrast to the quadratic complexity of attention mechanisms. These results demonstrate that SSM-based architectures are more effective and efficient for multi-task dense prediction tasks, where we need to process high-resolution feature maps in pixel-level prediction.
4.3.3 Effectiveness of Each Decoder Stage
As shown in Figure 1, the decoder of MTMamba++ consists of three stages. In this experiment, we study the effectiveness of these three stages on the NYUDv2 dataset. Table V presents the ablation results, clearly demonstrating that each decoder stage contributes positively to the performance of MTMamba++. The progressive performance gains achieved by successively incorporating each stage validate the effectiveness of our multi-stage decoder design in capturing and integrating multi-scale contextual features. As visualized in Figure 5, this hierarchical feature aggregation enables progressively refined predictions with sharper boundaries, particularly benefiting the boundary detection task.
TABLE IV: Comparison between SSM and attention on NYUDv2. We replace the SSM-related modules in MTMamba++ (i.e., the SS2D and CSS2D modules) with attention-based mechanisms (i.e., self-attention and cross-attention mechanisms).
@@ 4.3.3 Effectiveness of Each Decoder Stage
attention-based 55.15 0.4945 18.72 79.00 +2.63 448 796
SSM-based 57.01 0.4818 18.27 79.40 +4.82 315 524
@@
Figure 5: A qualitative comparison of each decoder stage in MTMamba++ on NYUDv2. Zoom in for more details.
@@
TABLE VII: Effect of expand factor $\alpha$ in MTMamba++ on NYUDv2 with different numbers of tasks. “S”, “D”, “N”, and “B” denote the semantic segmentation, depth estimation, surface normal estimation, and boundary detection tasks, respectively.
$\alpha$ Semseg Depth Normal Boundary $\Delta_{m}$[%]
@@
S-D 111 58.10 0.4768 - - +7.33
222 58.25 0.4808 - - +7.08
333 58.23 0.4859 - - +6.57
S-D-N 111 54.85 0.4956 18.57 - +2.79
222 55.78 0.4888 18.43 - +4.04
333 55.30 0.4932 18.39 - +3.53
S-D-N-B 111 56.74 0.4927 18.45 79.40 +3.93
222 57.01 0.4818 18.27 79.40 +4.82
333 55.85 0.4882 18.40 79.10 +3.71
@@ 4.3.4 Effect of Each Scan in CSS2D Module
4.3.5 Analysis of $\alpha$
As mentioned in Sections 3.5 and 3.6.2, in MTMamba++, both STM and S-CTM blocks expand the feature channel to improve the model capacity by a hyperparameter $\alpha$. We conduct an experiment on NYUDv2 to explore the relationship between $\alpha$ and task conflicts. Increasing the expansion factor $\alpha$ enhances the model’s representational capacity for capturing both task-specific features and cross-task interactions. However, excessively large values can lead to increased computational complexity and over-parameterization. The redundancy in the representation space dilutes effective information and makes model optimization more challenging, resulting in worse performance.
The results in Table VII demonstrate that the optimal $\alpha$ value is correlated with the severity of task conflicts. For the 2-task setting (S-D), $\alpha=1$ achieves the best $\Delta_{m}$, because the semantic segmentation and depth estimation tasks have relatively low conflict, requiring minimal additional capacity for cross-task interaction modeling. However, when the normal estimation task is added in the 3-task setting (S-D-N), task conflicts become more severe as evidenced by the significant performance drop of both semantic segmentation and depth estimation tasks. In this case, $\alpha=2$ becomes optimal in terms of $\Delta_{m}$, indicating that increased representational capacity is needed to handle the heightened task conflicts. In the 4-task setting (S-D-N-B), while the boundary detection task is added, the conflicts appear to be somewhat alleviated as the boundary detection task can provide complementary information to other tasks. Thus, $\alpha=2$ continues to perform best in terms of $\Delta_{m}$, maintaining the balance between adequate capacity for conflict resolution and avoiding over-parameterization. Notably, $\alpha=3$ consistently underperforms across all configurations, demonstrating that excessively large expansion factors lead to over-parameterization.
These results demonstrate that smaller $\alpha$ suffices for low-conflict scenarios, while moderately larger $\alpha$ is beneficial when severe conflicts exist, but excessively large $\alpha$ always degrades performance. Thus, $\alpha=2$ is adopted as the default configuration in MTMamba++ as it provides robust performance across various multi-task scenarios.
Figure 6: Qualitative comparison with baselines (i.e., InvPT [9], TaskPrompter [12], and MTMamba [24]) on the NYUDv2 dataset. As highlighted, MTMamba++ generates better predictions with more accurate details and sharper boundaries. In the semantic segmentation task, the black regions in GT denote the background and are excluded from the computation of loss and evaluation metric (i.e., mIoU). Zoom in for more details.
@@ 4.3.7 Analysis of Model Size and Cost
Table IX compares model size and FLOPs between the proposed MTMamba++ and baselines on the PASCAL-Context dataset. We can see that MTMamba++ achieves state-of-the-art performance while maintaining high computational efficiency. Specifically, with only 343MB parameters (14.3%, 18.7%, 15.7%, and 17.5% fewer than InvPT, TSP-Transformer, MLoRE, and TaskDiffusion, respectively), our MTMamba++ achieves superior performance across most tasks. In terms of computational cost, MTMamba++ requires only 609GB FLOPs, which is merely 30.6% of the resources needed by TSP-Transformer (1991GB) while still outperforming it. Compared to MLoRE and TaskDiffusion, MTMamba++ achieves better results with comparable computational demands. These results confirm that MTMamba++ offers not only performance advantages but also practical benefits for real-world applications through its efficient use of computational resources.
Figure 7: Qualitative comparison with baselines (i.e., InvPT [9], TaskPrompter [12], and MTMamba [24]) on the PASCAL-Context dataset. As highlighted, MTMamba++ generates better predictions with sharper boundaries and greater precision in small objects. Zoom in for more details.
Figure 8: Qualitative comparison with baselines (i.e., InvPT [9], TaskPrompter [12], and MTMamba [24]) on the Cityscapes dataset. As highlighted, MTMamba++ produces more precise predictions in small objects. Zoom in for more details.
@@ 4.4 Visualization of Predictions
In this section, we compare the output predictions from MTMamba++ against baselines, including InvPT [9], TaskPrompter [12], and MTMamba [24]. Figures 6, 7, and 8 show the qualitative results on NYUDv2, PASCAL-Context, and Cityscapes datasets, respectively. As can be seen, MTMamba++ has better visual results than baselines in all datasets. For example, as highlighted with yellow circles in Figure 6, MTMamba++ demonstrates fewer misclassification errors in semantic segmentation and produces sharper predicted boundaries in the boundary detection task. Figure 7 illustrates that MTMamba++ achieves more accurate detection of small objects in both semantic segmentation and human parsing tasks, particularly evident in the highlighted regions where our method can effectively detect distant pedestrians. MTMamba++ also generates sharper predicted boundaries for the object boundary detection task. Similarly, as highlighted in Figure 8, MTMamba++ achieves higher precision in detecting small objects (e.g., street lamps and tree trunks) in semantic segmentation, which are missed by Transformer-based methods. Hence, both qualitative study (Figures 6, 7, and 8) and quantitative study (Tables I and II) show the superior performance of MTMamba++.
@@ 5 Conclusion
In this paper, we propose MTMamba++, a novel multi-task architecture with a Mamba-based decoder for multi-task dense scene understanding. With two types of core blocks (i.e., STM and CTM blocks), MTMamba++ can effectively model long-range dependency and achieve cross-task interaction. We design two variants of the CTM block to promote knowledge exchange across tasks from the feature and semantic perspectives, respectively. Experiments on three benchmark datasets demonstrate that MTMamba++ achieves better performance than previous methods while maintaining high computational efficiency.